# Improving Low-Resource Knowledge Tracing Tasks by Supervised Pre-training and Importance Mechanism Fine-tuning

### v2

@@ 1 Introduction
Knowledge tracing holds a pivotal role within the realm of Intelligent Tutoring Systems (ITS) (Zhang and Yao, 2018; Su et al., 2021; Song et al., 2022; Ke et al., 2024). Its primary objective is to forecast students’ performance on questions by estimating their mastery of individual knowledge components (KCs [^1] A KC is a generality of everyday terms like concept, principle, or skill.) through an analysis of their past interactions. A KC is a description of a mental structure or process that a learner uses, alone or in combination with other KCs, to accomplish steps in a task or a problem. Take Figure 1 as an example. The student has successively responded to four questions ($Q_{1}$ to $Q_{4}$), achieving correct answers for $Q_{1}$ and $Q_{3}$, while $Q_{2}$ is answered incorrectly. This pattern suggests that the student may have a proficient understanding of the “Addition”, “Subtraction”, and “Multiplication” KCs, but lacks familiarity with the “Modulo” and “Division” KCs. Leveraging the current knowledge mastery, the KT task aims to predict the student’s performance on the upcoming sixth question, $Q_{4}$. After gaining insights into students’ knowledge mastery through KT, educators can promptly pinpoint weaknesses and provide targeted exercises for improvement. Additionally, this information can assist online learning platforms in providing a series of adaptive learning services such as learning resource recommendations, customizing student learning paths, and personalizing teaching plans (Liu et al., 2019; Wu et al., 2020).
@@
However, due to the users’ privacy protection of the educational applications and the different learning energy and enthusiasm of students, it is extremely difficult to collect large-scale high-quality student interaction sequences from real-world educational environments. Therefore, the educational datasets for KT model training frequently involve limited student learning records. For example, ASSISTments2009 is one of the classical KT datasets and the observed interaction records are only collected from 4,217 students. However, most of the state-of-the-art DLKT models are designed with stacks of neural networks such as recurrent neural networks, memory networks (Zhang et al., 2017; Nagatani et al., 2019; Sonkar et al., 2020; Guo et al., 2021). Directly training a DLKT model on such a low-resource KT dataset is very easy to run into the problem of overfitting. Furthermore, it is unclear what type of model architectures are most suitable for low-resource KT datasets in previous KT works.
To enhance the learning capability from low-resource datasets of deep learning-based models, some studies perform “pre-training and fine-tuning” paradigm (Zoph et al., 2016; Tu et al., 2019; Brown et al., 2020; Gira et al., 2022). This paradigm leverages rich-resource datasets to pre-train a model first and transfer the learned parameters to the low-resource dataset. Motivated by these promising studies, we propose a simple yet effective framework called LoReKT. LoReKT aims to improve the performance on the low-resource KT dataset by transferring the knowledge tracing capability from the model pre-trained on multiple rich-resource KT datasets. More specifically, in the pre-training stage, we build a foundational pre-trained KT model using a stack of transformer decoders based on multiple rich-resource KT datasets. To enhance the model’s capacity for integrating information from both questions and concepts, we introduce data type embeddings. Furthermore, to enable the model to learn the distinct and shared tracing patterns from multiple KT datasets, we introduce dedicated dataset embeddings for each KT dataset. In the fine-tuning stage, we propose an importance vector-based fine-tuning strategy to allow the model to focus on updating crucial parameters for the specific target low-resource dataset while constraining unimportant parameters to prevent the learning and memorization of noisy information.
@@
    To ensure that our approach can be fairly comparable with other recently developed DLKT models, we follow a publicly available standardized KT task evaluation protocol (Liu et al., 2022). We conduct comprehensive and rigorous experiments on three public rich-resource datasets and three low-resource datasets. The results show that after pre-training, the pre-trained KT model comes close to the performance of previous approaches on rich-resource datasets. By fine-tuning the pre-trained KT model on the low-resource datasets, it achieves superior prediction performance in terms of AUC and Accuracy compared to 17 baselines.
@@ 2.1 Deep Learning based Knowledge Tracing
Deep Knowledge Tracing (DKT) has pioneered the application of deep learning in knowledge tracing tasks by employing a Long Short Term Memory (LSTM) layer to encapsulate students’ knowledge states and predict students’ response performances (Piech et al., 2015). Since then, many methods tend to use deep learning techniques to solve KT problem (Zhang and Yao, 2018; Su et al., 2021; Song et al., 2022; Ke et al., 2024; Zhang et al., 2017; Yeung and Yeung, 2018; Nagatani et al., 2019; Lee and Yeung, 2019; Pandey and Karypis, 2019; Ghosh et al., 2020; Tiana et al., 2021; Song et al., 2022; Guo et al., 2021; Ma et al., 2022; Liu et al., 2023). For example, Yeung and Yeung (2018) leveraged prediction-consistent regularization mechanism to mitigate issues related to input reconstruction failure and prediction inconsistency in the context of DKT Piech et al. (2015). Zhang et al. (2017) integrated a meticulously designed static key matrix for storing the interconnections among different knowledge components (KCs). Simultaneously, it utilizes a dynamic value matrix to iteratively update the knowledge state of students. Motivated by the learning curve theory Murray et al. (2013)Nagatani et al. (2019) took student’s forgetting behavior into consideration to enhance DKT Piech et al. (2015). Lee and Yeung (2019) used student knowledge state encoder and skill encoder to predict the student response performance via the dot product. Tiana et al. (2021) performed multi-task learning based on the bidirectional encoder representations to construct mixed representations of questions. To mitigate the potential issue of limited generalization in DLKT, adversarial training techniques, such as adversarial perturbations, are introduced to the original student interaction sequence. Specifically, Guo et al. (2021) improveed the generalization capability of the DLKT model by incorporating adversarial perturbations at the embedding level of the student interaction sequence. The carefully designed perturbations contribute to the model’s effective generalization across diverse student interactions. Moreover, certain studies have concentrated on exploring the interactions between student responses and questions, as well as the associations between questions and KCs. For example, Nakagawa et al. (2019) constructed a question-concept knowledge graph and utilized graph neural network to aggregate the node features related to the corresponding concepts and subsequently updates the student’s knowledge states effectively. Additionally, Ma et al. (2022) employed self-supervised learning paradigm to identify the latent relationship between questions and KCs, thereby enhancing input representations. Another research direction focused on the interdependence among student interactions, aiming to capture finer details embedded within them. For instance, Pandey and Karypis (2019) utilized a self-attention mechanism to grasp the relationships between exercises and students’ responses. Ghosh et al. (2020) presented AKT, which utilizes two self-attention modules to extract the inner relevance of questions and interactions respectively, and explicitly model students’ forgetting behaviors via a monotonic attention mechanism.
@@ 2.2 Pre-training for Low-resource Setting
In real-world scenarios, encounters with low-resource settings are commonplace, and the paradigm of ”pre-training and fine-tuning” has consistently proven its efficacy in addressing challenges within such contexts (Yosinski et al., 2014; Shang et al., 2019; Bansal et al., 2019; Zhang et al., 2021; Liu et al., 2022; Chi et al., 2023; Zhang et al., 2023). For example, Yosinski et al. (2014) explored the transferability of AlexNet and observed that the initial three layers of AlexNet encapsulate general features conducive to transferability. By introducing fine-tuning to the neural network, it successfully mitigated data variability and scarcity, consequently enhancing the overall network performance. Bansal et al. (2019) introduced a straightforward methodology to enhance direct speech-to-text translation (ST) in scenarios where the source language is low-resource. The approach involves initial pre-training of the model on a high-resource automatic speech recognition (ASR) task, followed by a subsequent fine-tuning process to refine its parameters specifically for speech-to-text translation (ST). Zhang et al. (2021) proposed an innovative approach, the adaptive data augmentation fine-tuning technique, designed to facilitate the efficient transfer of Named Entity Recognition (NER) knowledge from resource-rich domains to low-resource target domains. Liu et al. (2022) utilized pre-training data in both the initial pre-training and subsequent fine-tuning stages, strategically enhancing the model’s performance across low-resource datasets. This dual-stage utilization of pre-training data contributes to a comprehensive and effective optimization, addressing the challenges posed by limited data availability in low-resource scenarios. Chi et al. (2023) investigated the prospect of enhancing the performance of low-resource non-English languages by incorporating pre-trained language models that are primarily trained on English. This exploration seeks to leverage the knowledge embedded in English-dominant language models to boost the capabilities of models applied to non-English languages with limited resources. To generate high-quality definition for low-resource language, Zhang et al. (2023) leveraged a multilingual pre-trained model as backbone and employed a prompt contrastive fine-tuning approach to enhance the model’s capabilities in this specific linguistic context.
@@ 4.1.1 Interaction Encoding
Due to the fact that the question bank is typically much larger than KCs, previous research mainly used KCs for interaction encoding, treating questions with the same KCs as identical (Piech et al., 2015; Yeung and Yeung, 2018; Nagatani et al., 2019). However, this approach led DLKT models to overlook the unique characteristics of same-KC questions, limiting interaction representation. To address this, we align with works that include both individual question features and KCs to encode interactions in a more granular manner (Choi et al., 2020; Ghosh et al., 2020; Liu et al., 2023).
@@
Drawing inspiration from the powerful pre-trained model BERT (Devlin et al., 2019), which leverages token type embeddings to improve the integration of various token information, we introduce data type embeddings to the KT problem. In this problem, there are two distinct data types: questions and concepts. To improve the pre-trained KT model’s ability to incorporate information from both, we introduce question and concept data type embeddings, which are directly integrated into all question and concept embeddings:
@@ 4.1.2 Pre-training
Recently, generative pre-trained models that based on Transformer architecture have achieved promising results in various tasks compared to designing a sophisticated neural network for a specific task (Radford et al., 2018; Brown et al., 2020; Gira et al., 2022). Drawing inspiration from these impressive findings, we opt to directly utilize a stack of transformer decoders. This choice enables us to dynamically capture student knowledge states without additional architecture design efforts:
@@
Furthermore, inspired by prompt learning techniques that effectively capture diverse and overlapping patterns in multi-task learning scenarios (Raffel et al., 2020; Lester et al., 2021; Sanh et al., 2022; Chung et al., 2022), we introduce dedicated dataset embeddings for individual KT datasets. Given that each KT dataset has its unique prediction paradigm stemming from variations in question banks and KCs, this enhancement empowers the pre-trained KT model to effectively capture the specific and shared information across different KT datasets. Specifically, the knowledge state $\mathbf{h}^{(l)}$ is first concatenated with corresponding dataset embedding $\mathbf{d}_{i}$ and QC embedding $\mathbf{x}_{t+1}$, then fed into a two-layer fully connected network with Sigmoid activation function $\sigma(\cdot)$ to predict the performance of a student on next question $q_{t+1}$:
@@ 4.2.1 Computing Importance Vector of Layer
The backbone of LoReKT is a stack of transformer decoders. The key components of a transformer decoder are multi-head attention layer, intermediate layer, and output layer [^5] In this paper, we use “layer” or $l$ to indicate any of these three layers, because the procedure of computing these three layers’ importance vector is similar.. It has been found that not all units (neurons or attention heads) in a specific layer are important (Michel et al., 2019). Therefore, before directly fine-tuning the model on each low-resource KT dataset, we adopt the approach described by Ke et al. (2022) to compute the importance vector for each layer. This is achieved by employing a gradient-based importance detection method, which is specifically tailored for each low-resource KT dataset:
@@ 5.1 Datasets
Since our LoReKT framework learns transferable parameters and representations from rich-resource KT datasets first and then quickly adapt to low-resource scenarios, we select three rich-resource datasets including BD2006 (Stamper et al., 2010), XES3G5M (Liu et al., 2023), and EdNet (Choi et al., 2020) to establish a robust foundational pre-trained KT model. We further fine-tune the pre-trained KT model on three low-resource KT datasets including AS2009 (Feng et al., 2009), NIPS34 (Wang et al., 2020), and AL2005 (Stamper et al., 2010) respectively. The data statistics for the six selected datasets can be found in Table 1. The detailed descriptions are as follows:
@@ 5.2 Experimental Setting
We remove student sequences shorter than 3 attempts and truncate student interaction sequences that are longer than 200. We use 80% of student sequences for training and validation and the rest 20% of student sequences for model evaluation. We adopt Adam optimizer (Kingma and Ba, 2015) to train all the models. The number of training epochs is set to 200. Following all existing DLKT research (Liu et al., 2022, 2023; Choi et al., 2020), we use the Area Under the Curve (AUC) as the main evaluation metric and use Accuracy as the secondary evaluation metric. We use an early stopping strategy that stops optimization when the AUC score fails to get the improvement on the validation set in the latest 10 epochs. Owing to the training efficiency, it is difficult to tune too many hyperparameters for the models with billions of parameters. Hence, we only tune the model learning rate in {0.001, 0.0001} with the dropout rate in {0.1, 0.2} for fair comparison in various model sizes.
@@ 5.3 Baselines
    DKT Piech et al. (2015): it uses an LSTM layer to model students’ learning processes.
@@
    DKT+ Yeung and Yeung (2018): it improves the original DKT model by addressing the reconstruction and inconsistent issues.
@@
    DKT-F Nagatani et al. (2019): it enhances original DKT by considering students’ forgetting behaviors.
@@
    KQN Lee and Yeung (2019): it is a recurrent neural network (RNN) based architecture that extracts the relation representations between students’ learning abilities and KCs to predict their performance.
@@
    LPKT Shen et al. (2021): it designs a learning cell to model the students’ learning processes to estimate their knowledge states.
@@
    AT-DKT Liu et al. (2023): it proposes two auxiliary learning tasks involving question tagging prediction task and individualized prior knowledge prediction task to improve the prediction performance of DKT.
@@
    DKVMN Zhang et al. (2017): it incorporates a static matrix to store the relationships among KCs and a dynamic matrix to track the student’s knowledge state.
@@
    SKVMN Abdelrahman and Wang (2019): it is a combination of DKVMN and LSTM that uses a hop-LSTM layer to capture sequential dependencies of questions.
@@
    DeepIRT Yeung (2019): it incorporates DKVMN and item response theory to enhance the interpretability of the prediction output of DKVMN.
@@
    SAKT Pandey and Karypis (2019): it uses self-attention to identify the relevance between historical interactions and KCs.
@@
    SAINT Choi et al. (2020): it is a Transformer-based model for KT that encodes questions and responses in the encoder and decoder respectively.
@@
    AKT Ghosh et al. (2020): it leverages three self-attention modules to estimate the relevance between questions and historical interactions and explicitly models student’s forgetting behavior via a monotonic attention mechanism.
@@
    simpleKT Liu et al. (2023): it explores the ordinary dot-product attention based KT models by capturing the individual differences among questions covering the same set of KCs
@@
    sparseKT Huang et al. (2023): it incorporates a k-selection module to only pick items with the highest attention scores to improve the robustness and generalization of the attention based DLKT approaches.
@@
    HawkesKT Wang et al. (2021): it utilizes the Hawkes process to model temporal cross-effects in student historical interactions.
@@
    ATKT Guo et al. (2021): it performs adversarial perturbations into the student interaction sequence to enhance the generalization ability based on an attention-LSTM based KT model.
@@
    GKT Nakagawa et al. (2019): it casts the knowledge structure as a graph and reformulates the KT task as a time series node-level classification problem in GNN.
@@
DKT (Piech et al., 2015) Sequential 0.7525 0.7688 0.8159 0.8018 0.7845 0.6405
DKVMN (Zhang et al., 2017) Memory 0.7472 0.7677 0.8052 0.7999 0.7796 0.6576
DKT+ (Yeung and Yeung, 2018) Sequential 0.7543 0.7698 0.8141 0.8019 0.7858 0.6454
DKT-F (Nagatani et al., 2019) Sequential - 0.7728 0.8146 0.7997 0.7935 0.6548
KQN (Lee and Yeung, 2019) Sequential 0.7462 0.7685 0.8010 0.7953 0.7794 0.6415
SKVMN (Abdelrahman and Wang, 2019) Memory 0.7332 0.7513 0.7463 0.7287 0.7512 0.6374
DeepIRT (Yeung, 2019) Memory 0.7465 0.7673 0.8040 0.7976 0.7789 0.6387
GKT (Nakagawa et al., 2019) Others 0.7442 0.7718 0.8112 0.8041 0.7731 0.6392
SAKT (Pandey and Karypis, 2019) Attention 0.7221 0.7508 0.7850 0.7748 0.7685 0.6290
SAINT (Choi et al., 2020) Attention 0.6990 0.7883 0.7764 0.7758 0.8070 0.6841
AKT (Ghosh et al., 2020) Attention 0.7869 0.8038 0.8324 0.8213 0.8215 0.7054
ATKT (Guo et al., 2021) Others 0.7472 0.7664 0.7987 0.7889 0.7791 0.6490
HawkesKT (Wang et al., 2021) Others 0.7232 0.7763 0.8199 0.8077 0.7933 0.7304
LPKT (Shen et al., 2021) Sequential 0.7812 0.8004 0.8268 0.8056 0.8163 0.7644
AT-DKT (Liu et al., 2023) Sequential 0.7555 0.7816 0.8246 0.8104 0.7925 0.6536
simpleKT (Liu et al., 2023) Attention 0.7744 0.8035 0.8254 0.8160 0.8161 0.6765
sparseKT (Huang et al., 2023) Attention 0.7739 0.8033 0.8152 0.8120 0.8165 0.6804

### v3

@@ 1 Introduction
Knowledge tracing holds a pivotal role within the realm of Intelligent Tutoring Systems (ITS) [1, 2, 3, 4]. Its primary objective is to forecast students’ performance on questions by estimating their mastery of individual knowledge components (KCs [^1] A KC is a generality of everyday terms like concept, principle, or skill.) through an analysis of their past interactions. A KC is a description of a mental structure or process that a learner uses, alone or in combination with other KCs, to accomplish steps in a task or a problem. Take Figure 1 as an example. The student has successively responded to four questions ($Q_{1}$ to $Q_{4}$), achieving correct answers for $Q_{1}$ and $Q_{3}$, while $Q_{2}$ is answered incorrectly. This pattern suggests that the student may have a proficient understanding of the “Addition”, “Subtraction”, and “Multiplication” KCs, but lacks familiarity with the “Modulo” and “Division” KCs. Leveraging the current knowledge mastery, the KT task aims to predict the student’s performance on the upcoming sixth question, $Q_{4}$. After gaining insights into students’ knowledge mastery through KT, educators can promptly pinpoint weaknesses and provide targeted exercises for improvement. Additionally, this information can assist online learning platforms in providing a series of adaptive learning services such as learning resource recommendations, customizing student learning paths, and personalizing teaching plans [5, 6].
@@
However, due to the users’ privacy protection of the educational applications and the different learning energy and enthusiasm of students, it is extremely difficult to collect large-scale high-quality student interaction sequences from real-world educational environments. Therefore, the educational datasets for KT model training frequently involve limited student learning records. For example, ASSISTments2009 is one of the classical KT datasets and the observed interaction records are only collected from 4,217 students. However, most of the state-of-the-art DLKT models are designed with stacks of neural networks such as recurrent neural networks, memory networks [7, 8, 9, 10]. Directly training a DLKT model on such a low-resource KT dataset is very easy to run into the problem of overfitting. Furthermore, it is unclear what type of model architectures are most suitable for low-resource KT datasets in previous KT works.
To enhance the learning capability from low-resource datasets of deep learning-based models, some studies perform “pre-training and fine-tuning” paradigm [11, 12, 13, 14]. This paradigm leverages rich-resource datasets to pre-train a model first and transfer the learned parameters to the low-resource dataset. Motivated by these promising studies, we propose a simple yet effective framework called LoReKT. LoReKT aims to improve the performance on the low-resource KT dataset by transferring the knowledge tracing capability from the model pre-trained on multiple rich-resource KT datasets. More specifically, in the pre-training stage, we build a foundational pre-trained KT model using a stack of transformer decoders based on multiple rich-resource KT datasets. To enhance the model’s capacity for integrating information from both questions and concepts, we introduce data type embeddings. Furthermore, to enable the model to learn the distinct and shared tracing patterns from multiple KT datasets, we introduce dedicated dataset embeddings for each KT dataset. In the fine-tuning stage, we propose an importance vector-based fine-tuning strategy to allow the model to focus on updating crucial parameters for the specific target low-resource dataset while constraining unimportant parameters to prevent the learning and memorization of noisy information.
@@
    To ensure that our approach can be fairly comparable with other recently developed DLKT models, we follow a publicly available standardized KT task evaluation protocol [15]. We conduct comprehensive and rigorous experiments on three public rich-resource datasets and three low-resource datasets. The results show that after pre-training, the pre-trained KT model comes close to the performance of previous approaches on rich-resource datasets. By fine-tuning the pre-trained KT model on the low-resource datasets, it achieves superior prediction performance in terms of AUC and Accuracy compared to 17 baselines.
@@ 2.1 Deep Learning based Knowledge Tracing
Deep Knowledge Tracing (DKT) has pioneered the application of deep learning in knowledge tracing tasks by employing a Long Short Term Memory (LSTM) layer to encapsulate students’ knowledge states and predict students’ response performances [16]. Since then, many methods tend to use deep learning techniques to solve KT problem [1, 2, 3, 4, 7, 17, 8, 18, 19, 20, 21, 22, 10, 23, 24]. For example, Yeung and Yeung [17] leveraged prediction-consistent regularization mechanism to mitigate issues related to input reconstruction failure and prediction inconsistency in the context of DKT [16]. Zhang et al. [7] integrated a meticulously designed static key matrix for storing the interconnections among different knowledge components (KCs). Simultaneously, it utilizes a dynamic value matrix to iteratively update the knowledge state of students. Motivated by the learning curve theory [25]Nagatani et al. [8] took student’s forgetting behavior into consideration to enhance DKT [16]. Lee and Yeung [18] used student knowledge state encoder and skill encoder to predict the student response performance via the dot product. Tiana et al. [21] performed multi-task learning based on the bidirectional encoder representations to construct mixed representations of questions. To mitigate the potential issue of limited generalization in DLKT, adversarial training techniques, such as adversarial perturbations, are introduced to the original student interaction sequence. Specifically, Guo et al. [10] improveed the generalization capability of the DLKT model by incorporating adversarial perturbations at the embedding level of the student interaction sequence. The carefully designed perturbations contribute to the model’s effective generalization across diverse student interactions. Moreover, certain studies have concentrated on exploring the interactions between student responses and questions, as well as the associations between questions and KCs. For example, Nakagawa et al. [26] constructed a question-concept knowledge graph and utilized graph neural network to aggregate the node features related to the corresponding concepts and subsequently updates the student’s knowledge states effectively. Additionally, Ma et al. [23] employed self-supervised learning paradigm to identify the latent relationship between questions and KCs, thereby enhancing input representations. Another research direction focused on the interdependence among student interactions, aiming to capture finer details embedded within them. For instance, Pandey and Karypis [19] utilized a self-attention mechanism to grasp the relationships between exercises and students’ responses. Ghosh et al. [20] presented AKT, which utilizes two self-attention modules to extract the inner relevance of questions and interactions respectively, and explicitly model students’ forgetting behaviors via a monotonic attention mechanism.
@@ 2.2 Pre-training for Low-resource Setting
In real-world scenarios, encounters with low-resource settings are commonplace, and the paradigm of ”pre-training and fine-tuning” has consistently proven its efficacy in addressing challenges within such contexts [27, 28, 29, 30, 31, 32, 33]. For example, Yosinski et al. [27] explored the transferability of AlexNet and observed that the initial three layers of AlexNet encapsulate general features conducive to transferability. By introducing fine-tuning to the neural network, it successfully mitigated data variability and scarcity, consequently enhancing the overall network performance. Bansal et al. [29] introduced a straightforward methodology to enhance direct speech-to-text translation (ST) in scenarios where the source language is low-resource. The approach involves initial pre-training of the model on a high-resource automatic speech recognition (ASR) task, followed by a subsequent fine-tuning process to refine its parameters specifically for speech-to-text translation (ST). Zhang et al. [30] proposed an innovative approach, the adaptive data augmentation fine-tuning technique, designed to facilitate the efficient transfer of Named Entity Recognition (NER) knowledge from resource-rich domains to low-resource target domains. Liu et al. [31] utilized pre-training data in both the initial pre-training and subsequent fine-tuning stages, strategically enhancing the model’s performance across low-resource datasets. This dual-stage utilization of pre-training data contributes to a comprehensive and effective optimization, addressing the challenges posed by limited data availability in low-resource scenarios. Chi et al. [32] investigated the prospect of enhancing the performance of low-resource non-English languages by incorporating pre-trained language models that are primarily trained on English. This exploration seeks to leverage the knowledge embedded in English-dominant language models to boost the capabilities of models applied to non-English languages with limited resources. To generate high-quality definition for low-resource language, Zhang et al. [33] leveraged a multilingual pre-trained model as backbone and employed a prompt contrastive fine-tuning approach to enhance the model’s capabilities in this specific linguistic context.
@@ 4.1.1 Interaction Encoding
Due to the fact that the question bank is typically much larger than KCs, previous research mainly used KCs for interaction encoding, treating questions with the same KCs as identical [16, 34, 35]. However, this approach led DLKT models to overlook the unique characteristics of same-KC questions, limiting interaction representation. To address this, we align with works that include both individual question features and KCs to encode interactions in a more granular manner [36, 20, 24].
@@
Drawing inspiration from the powerful pre-trained model BERT [37], which leverages token type embeddings to improve the integration of various token information, we introduce data type embeddings to the KT problem. In this problem, there are two distinct data types: questions and concepts. To improve the pre-trained KT model’s ability to incorporate information from both, we introduce question and concept data type embeddings, which are directly integrated into all question and concept embeddings:
@@ 4.1.2 Pre-training
Recently, generative pre-trained models that based on Transformer architecture have achieved promising results in various tasks compared to designing a sophisticated neural network for a specific task [38, 13, 14]. Drawing inspiration from these impressive findings, we opt to directly utilize a stack of transformer decoders. This choice enables us to dynamically capture student knowledge states without additional architecture design efforts:
@@
Furthermore, inspired by prompt learning techniques that effectively capture diverse and overlapping patterns in multi-task learning scenarios [39, 40, 41, 42], we introduce dedicated dataset embeddings for individual KT datasets. Given that each KT dataset has its unique prediction paradigm stemming from variations in question banks and KCs, this enhancement empowers the pre-trained KT model to effectively capture the specific and shared information across different KT datasets. Specifically, the knowledge state $\mathbf{h}^{(l)}$ is first concatenated with corresponding dataset embedding $\mathbf{d}_{i}$ and QC embedding $\mathbf{x}_{t+1}$, then fed into a two-layer fully connected network with Sigmoid activation function $\sigma(\cdot)$ to predict the performance of a student on next question $q_{t+1}$:
@@ 4.2.1 Computing Importance Vector of Layer
The backbone of LoReKT is a stack of transformer decoders. The key components of a transformer decoder are multi-head attention layer, intermediate layer, and output layer [^5] In this paper, we use “layer” or $l$ to indicate any of these three layers, because the procedure of computing these three layers’ importance vector is similar.. It has been found that not all units (neurons or attention heads) in a specific layer are important [43]. Therefore, before directly fine-tuning the model on each low-resource KT dataset, we adopt the approach described by Ke et al. [44] to compute the importance vector for each layer. This is achieved by employing a gradient-based importance detection method, which is specifically tailored for each low-resource KT dataset:
@@ 5.1 Datasets
Since our LoReKT framework learns transferable parameters and representations from rich-resource KT datasets first and then quickly adapt to low-resource scenarios, we select three rich-resource datasets including BD2006 [47], XES3G5M [48], and EdNet [49] to establish a robust foundational pre-trained KT model. We further fine-tune the pre-trained KT model on three low-resource KT datasets including AS2009 [50], NIPS34 [51], and AL2005 [47] respectively. The data statistics for the six selected datasets can be found in Table 1. The detailed descriptions are as follows:
@@ 5.2 Experimental Setting
We remove student sequences shorter than 3 attempts and truncate student interaction sequences that are longer than 200. We use 80% of student sequences for training and validation and the rest 20% of student sequences for model evaluation. We adopt Adam optimizer [52] to train all the models. The number of training epochs is set to 200. Following all existing DLKT research [15, 24, 36], we use the Area Under the Curve (AUC) as the main evaluation metric and use Accuracy as the secondary evaluation metric. We use an early stopping strategy that stops optimization when the AUC score fails to get the improvement on the validation set in the latest 10 epochs. Owing to the training efficiency, it is difficult to tune too many hyperparameters for the models with billions of parameters. Hence, we only tune the model learning rate in {0.001, 0.0001} with the dropout rate in {0.1, 0.2} for fair comparison in various model sizes.
@@ 5.3 Baselines
    DKT [16]: it uses an LSTM layer to model students’ learning processes.
@@
    DKT+ [17]: it improves the original DKT model by addressing the reconstruction and inconsistent issues.
@@
    DKT-F [8]: it enhances original DKT by considering students’ forgetting behaviors.
@@
    KQN [18]: it is a recurrent neural network (RNN) based architecture that extracts the relation representations between students’ learning abilities and KCs to predict their performance.
@@
    LPKT [53]: it designs a learning cell to model the students’ learning processes to estimate their knowledge states.
@@
    AT-DKT [54]: it proposes two auxiliary learning tasks involving question tagging prediction task and individualized prior knowledge prediction task to improve the prediction performance of DKT.
@@
    DKVMN [7]: it incorporates a static matrix to store the relationships among KCs and a dynamic matrix to track the student’s knowledge state.
@@
    SKVMN [55]: it is a combination of DKVMN and LSTM that uses a hop-LSTM layer to capture sequential dependencies of questions.
@@
    DeepIRT [56]: it incorporates DKVMN and item response theory to enhance the interpretability of the prediction output of DKVMN.
@@
    SAKT [19]: it uses self-attention to identify the relevance between historical interactions and KCs.
@@
    SAINT [36]: it is a Transformer-based model for KT that encodes questions and responses in the encoder and decoder respectively.
@@
    AKT [20]: it leverages three self-attention modules to estimate the relevance between questions and historical interactions and explicitly models student’s forgetting behavior via a monotonic attention mechanism.
@@
    simpleKT [24]: it explores the ordinary dot-product attention based KT models by capturing the individual differences among questions covering the same set of KCs
@@
    sparseKT [57]: it incorporates a k-selection module to only pick items with the highest attention scores to improve the robustness and generalization of the attention based DLKT approaches.
@@
    HawkesKT [58]: it utilizes the Hawkes process to model temporal cross-effects in student historical interactions.
@@
    ATKT [10]: it performs adversarial perturbations into the student interaction sequence to enhance the generalization ability based on an attention-LSTM based KT model.
@@
    GKT [26]: it casts the knowledge structure as a graph and reformulates the KT task as a time series node-level classification problem in GNN.
@@
DKT [16] Sequential 0.7525 0.7688 0.8159 0.8018 0.7845 0.6405
DKVMN [7] Memory 0.7472 0.7677 0.8052 0.7999 0.7796 0.6576
DKT+ [17] Sequential 0.7543 0.7698 0.8141 0.8019 0.7858 0.6454
DKT-F [8] Sequential - 0.7728 0.8146 0.7997 0.7935 0.6548
KQN [18] Sequential 0.7462 0.7685 0.8010 0.7953 0.7794 0.6415
SKVMN [55] Memory 0.7332 0.7513 0.7463 0.7287 0.7512 0.6374
DeepIRT [56] Memory 0.7465 0.7673 0.8040 0.7976 0.7789 0.6387
GKT [26] Others 0.7442 0.7718 0.8112 0.8041 0.7731 0.6392
SAKT [19] Attention 0.7221 0.7508 0.7850 0.7748 0.7685 0.6290
SAINT [36] Attention 0.6990 0.7883 0.7764 0.7758 0.8070 0.6841
AKT [20] Attention 0.7869 0.8038 0.8324 0.8213 0.8215 0.7054
ATKT [10] Others 0.7472 0.7664 0.7987 0.7889 0.7791 0.6490
HawkesKT [58] Others 0.7232 0.7763 0.8199 0.8077 0.7933 0.7304
LPKT [53] Sequential 0.7812 0.8004 0.8268 0.8056 0.8163 0.7644
AT-DKT [54] Sequential 0.7555 0.7816 0.8246 0.8104 0.7925 0.6536
simpleKT [24] Attention 0.7744 0.8035 0.8254 0.8160 0.8161 0.6765
sparseKT [57] Attention 0.7739 0.8033 0.8152 0.8120 0.8165 0.6804